# Perturbation-based Analysis of Compositional Data

### v1

@@ 1 Introduction
The remainder of the paper is structured as follows. In Section 1.1 we discuss existing methods for compositional data analysis and how our approach contributes to this research. Section 1.2 collects the mathematical notation used throughout this work. Our perturbation-based framework is introduced in Section 2. In Section 2.1, we provide a way to construct interpretable perturbations based on suitably choosing the endpoint, direction and speed of the perturbation. We also provide a general construction for the aforementioned $L$ and $W$ for a wide range of perturbations. We end the section by discussing how to adapt our proposals to general geometries on the simplex including those employed by many current methods in compositional data analysis. In Section 2.2 we highlight and name some of the perturbation effects that we believe are most useful for practitioners and provide some examples of how and when these effects should be applied. In Section 3, we discuss how to estimate the parameters of interest based on semiparametric theory using two different methods that represent a trade-off between generality and robustness. Our semiparametric estimators and perturbation effects are then illustrated on both simulated data and two real-world applications in Section 4. Finally, we present the theoretical guarantees of our estimates in Section 5.
1.1 Relation to existing compositional data analysis approaches
Compositional data analysis has a long history dating back to Pearson (1897). The most prevalent class of methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach. In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space. By considering ratios instead of compositional coordinates it avoids the common pitfalls of spurious correlations that occur by comparing coordinates. Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3). Furthermore, some of the perturbation effects that we highlight as interpretable overlap with effects that can also be deduced from the log-ratio perspective. Most prominently, we consider a class of compositional feature influence measures for which a certain member (multiplicative speed, see Example 3) corresponds to the compositional feature influence which has been previously proposed by Huang et al. (2023). It also corresponds to one of the coefficients in a log-contrast model (Aitchison and Bacon-Shone, 1984) if the conditional mean lies in the log-contrast model class. Our perturbation-based framework, however, has a fundamentally different starting point and allows for the construction of perturbations that are not related to the log-ratio approach.
@@ 1.1 Relation to existing compositional data analysis approaches
The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however we believe its indirect use of the data is restrictive. This is best illustrated by the so-called zero-problem that occurs because the log-ratio approach, in general, cannot be applied in settings with too many zeros. This problem has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach. Other approaches advocate the use of different transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022). While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa. Greenacre et al. (2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches. In contrast to these methods, our perturbation-based framework works directly with the compositional data and defines the target parameter from this perspective. In the presence of zeros, perturbations that directly incorporate or omit zeros can be selected based on the application. Furthermore, we can construct perturbation effects that directly capture the effect of setting certain components to zero.
@@ 2 Target parameters via perturbations
    $v_{\psi}(z):=\mathbbm{1}_{\{\omega_{\psi}(z)=0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
@@
Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex(see Section S1 for a precise definition) at $z$, we define the average (directional) perturbation effect by
@@
We showcase different potential applications for both types of perturbation effect in the numerical experiments in Section 4. For binary perturbations, the denominator is included to scale the effect by the probability that the perturbation does not change $Z$. This is not required but in many cases provides a more meaningful parameter if one is interested in how the perturbation affects points on the simplex that are not yet at their endpoint. While both target parameters have obvious similarities to causal effects, they are a priori only (observational) statistical quantities. In particular, they are identifiable from the observational distribution as long as the conditions in Definition 2 are satisfied. We discuss the assumptions required for causal conclusions in Section 2.3.
Direct estimation of both target parameters relies on estimating the regression function $f:z\mapsto\mathbb{E}[Y\mid Z=z]$ and taking averages (see Section 3). An alternative estimation procedure for the effect of a binary perturbation $\lambda_{\psi}$ can be derived from noting that, if $L=\mathbbm{1}_{\{Z=\psi(Z,1)\}}$ and $W=\psi(Z,1)$, then
$\mathbb{E}[Y\,|\,L=1,W]=\frac{\mathbb{E}[YL\,|\,W]}{\mathbb{E}[L\,|\,W]}=f(W).$
@@ 2.1 Perturbations on the simplex
We now turn to the construction of perturbations on the simplex. For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1). For a fixed endpoint there is a canonical (potentially geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint. The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation. We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from a summary statistic and assuming it satisfies a partially linear model as in (5) with $L$ being the summary statistic. For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
@@ 2.1.1 Endpoints in the simplex
As formally defined in Definition 1, an endpoint function $E:\Delta^{d-1}\rightarrow\Delta^{d-1}$, if it exists, determines where the perturbation ends when $\gamma$ is increased as much as possible. The simplest type of endpoint function on the simplex is a constant function, that is, there exists a fixed point $z_{*}\in\Delta^{d-1}$ such that for all $z\in\Delta^{d-1}$ it holds that $E(z)=z_{*}$. For example, we can choose the fixed endpoint $z_{*}=e_{j}$, which corresponds to a vertex on the simplex and hence a perturbation with this endpoint moves proportions from the other coordinate into the $j$th one. A further example, is to use the center of the simplex, i.e. $z_{*}=z_{\mathrm{cen}}$. As alluded to in the introduction, perturbations with $z_{\mathrm{cen}}$ as the endpoint can be seen as perturbing observations in the direction of equal distribution and hence capture a notion of diversity. For later reference, we call all perturbation $\psi$ with endpoint function $E_{\psi}:z\mapsto z_{\mathrm{cen}}$ centering perturbations.
A second type of endpoint function, that can be non-constant, is given by mapping each $z\in\Delta^{d-1}$ to amalgamations on the simplex. An amalgamation is a well-studied operation (e.g., Aitchison, 2003) which consist of merging parts of the compositional components into others. Formally, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint. Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can be easily extended to settings where zero speeds are allowed. In Section 2.1.3 we describe how working in other geometries can lead to different notions of direction. Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect. We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, for which a partially linear model as in (5) can be assumed, and then constructing a corresponding speed. In both cases, the construction of a derivative-isolating reparametrization is similar.
@@
Intuitively, a derivative-isolating reparametrization for $\psi$ should isolate changes along the direction $v_{E}$ from changes due to varying directions. We therefore consider the reparametrization $\phi_{\mathrm{unit}}:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(\|E(z)-z\|_{1},(E(z),v_{E}(z)))$, where $\mathcal{W}$ is formally defined in Proposition 2. The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=-\ell w_{v}+w_{E}$, so
@@
and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$. Since these type of unit-speed perturbations may not always be easy to interpret, we extend this basic construction to start from either a given interpretable speed or a summary statistic. The proof of the following result is given in Section S2.2 of the supplementary material.
@@
For $\mathcal{Z}\subseteq\Delta^{d-1}$, let $E:\mathcal{Z}\to\Delta^{d-1}$ be an endpoint function satisfying for all $z\in\mathcal{Z}$ that $E(z)\neq z$. Denote by $\mathcal{E}:=E(\mathcal{Z})$ the set of endpoints and define for all $z^{*}\in\mathcal{E}$ and all $u\in\mathbb{R}^{d}$ the sets $\mathcal{I}_{(z^{*},u)}:=\{\delta\in(0,\infty)\mid z^{*}+\delta u\in\Delta^{d-1}\}$ and $\mathcal{W}:=\{(w_{E},w_{v})\in\mathcal{E}\times\mathbb{R}^{d}\mid\|w_{v}\|_{1}=1,\,\mathcal{I}_{w}\neq\emptyset\}$. Finally, define the unit speed reparametrization $\phi_{\mathrm{unit}}=(\phi_{\mathrm{unit}}^{L},\phi_{\mathrm{unit}}^{W}):\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}^{L}(z):=\|E(z)-z\|_{1}$ and $\phi_{\mathrm{unit}}^{W}(z):=(v(z),E(z))$ with $v$ the canonical endpoint direction defined in (6). Then the following approaches lead to perturbations with corresponding derivative-isolating reparametrizations.
@@
    (Speed) Let $s:\mathcal{Z}\rightarrow(0,\infty)$ be a given speed function. For all $w\in\mathcal{W}$ define $s_{w}:\mathcal{I}_{w}\rightarrow(0,\infty)$ for all $\delta\in\mathcal{I}_{w}$ by $s_{w}(\delta):=s(\phi_{\mathrm{unit}}^{-1}(\delta,w))$ and let $t_{w}:\mathcal{I}_{w}\to\mathbb{R}$ denote a solution to
@@
    Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t_{\phi_{\mathrm{unit}}^{W}(z)}(\|E(z)-z\|_{1}),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v(z)$ and any other perturbation with the same speed and direction.
@@
    (Summary statistic) Let $t:\mathcal{Z}\rightarrow\mathbb{R}$ be a given summary statistic. For all $w\in\mathcal{W}$ define $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ for all $\delta\in\mathcal{I}_{w}$ by $t_{w}(\delta):=t(\phi_{\mathrm{unit}}^{-1}(\delta,w))$ and assume that $t_{w}$ is differentiable and strictly decreasing. Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t(z),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v(z)$, where the speed $s:\mathcal{Z}\rightarrow\mathbb{R}$ is defined for all $z\in\mathcal{Z}$ as
@@
Suppose we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in\Delta^{d-1}$. We can investigate the effect of such a change by means of a $([d]\setminus\{j\}\to\{j\})$-amalgamating perturbation, that is, a perturbation which satisfies for all $z\in\Delta^{d-1}\setminus\{e_{j}\}$ that $v_{\psi}(z)=\frac{e_{j}-z}{\|e_{j}-z\|_{1}}$. While for some purposes it may be perfectly reasonable to use a constant speed $s_{\psi}\equiv 1$, in other settings this speed can be unintuitive. For example, if we imagine the observed $Z$ as being generated from an unobserved $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale by $Z:=\mathsf{C}(X)=\frac{1}{\|X\|_{1}}X$, it might be more natural to consider which speeds are induced on the simplex from hypothetical changes to $X^{j}$. A first attempt might be to consider the speed resulting from adding a constant $c$ to $X^{j}$, however, since
@@
this speed is not scale-invariant for $X$, that is, the speed when considering the counts $X$ and $cX$ for $c>0$ may differ. This means that the speed on the simplex would depend on the unobserved quantity $\|X\|_{1}$ so the perturbation is ill-defined. Instead, we can consider the speed resulting from multiplying $X^{j}$ by $1+c$ (so that $c=0$ corresponds to no change) which yields a well-defined perturbation on the simplex since this operation is scale-invariant. More formally, letting $\odot$ denote the Hadamard (point-wise) product, the resulting speed is
@@
which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition 2 (a). Since for all $z\in\mathcal{Z}$ it holds that $E(z)^{j}=1$ and $v(z)^{j}=\frac{1}{2}$, we have that $s_{w}$ (defined in Proposition 2) does not depend on $w$ and is of the form
$s_{w}:\delta\mapsto\frac{(2-\delta)\delta}{2}.$
@@
$t_{w}(\delta)=\log\left(\frac{z^{j}}{1-z^{j}}\right).$

### v2

@@ 1 Introduction
The remainder of the paper is structured as follows. In Section 1.1 we discuss existing methods for compositional data analysis and how our approach contributes to this research. Section 1.2 collects the mathematical notation used throughout this work. Our perturbation-based framework is introduced in Section 2. In Section 2.1, we provide a way to construct interpretable perturbations based on suitably choosing the endpoint, direction and speed of the perturbation. We also provide a general construction for the aforementioned $L$ and $W$ for a wide range of perturbations. We end the section by discussing how to adapt our proposals to general geometries on the simplex including those employed by many current methods in compositional data analysis. In Section 2.2 we highlight and name some of the perturbation effects that we believe are most useful for practitioners and provide some examples of how and when these effects should be applied. We provide a short discussion on how to connect the average perturbation effects to causal quantities in Section 2.3. In Section 3, we discuss how to estimate the parameters of interest based on semiparametric theory using two different methods that represent a trade-off between generality and robustness. Our semiparametric estimators and perturbation effects are then illustrated on both simulated and semi-synthetic data and applied to two real-world applications in Section 4. Finally, we present the theoretical guarantees of our estimates in Section 5.
1.1 Related literature
This paper addresses problems from several different fields of statistics. Our proposed perturbation-based framework is perhaps most closely related to the literature on causal inference, semiparametric inference and targeted learning (Bickel et al., 1993; Tsiatis, 2006; van der Laan and Rose, 2011). Spirtes and Scheines (2004) discuss some of the problems of modelling causal effects when certain variables are deterministic functions of other variables. This discussion was later continued by Woodward (2015) and new causal models have been developed to address this problem (Beckers et al., 2023). We contribute to these developments by providing a novel framework for causal effect estimation when some covariates are compositional (see Section 2.3).
An important aspect of causal effect estimation more generally is the need for an interpretable summary measure of the effect. Historically, such interpretable measures originated from parametric models that were easy to interpret. Much work has since gone into the interpretation of and uncertainty quantification for such parameters when the model is misspecified. Perhaps one of the most impactful such analyses was given by White (1980) who provided an estimation procedure for the coefficients in a linear model that did not require the model to be correctly specified. To what extent parameters based on linear models are useful and interpretable when the model is misspecified remains an active research topic particularly in econometrics (Aronow and Samii, 2016; Graham and de Xavier Pinto, 2022; Słoczyński, 2022). More recently, interpretable parameters based on partially linear models have been proposed and analyzed extensively (Chernozhukov et al., 2018; Vansteelandt and Dukes, 2022). Despite these and similar parameters depending on potentially complicated nuisance functions, it is often possible to estimate them efficiently under relatively mild conditions, as we also do here, by utilizing semiparametric theory (Kennedy, 2023).
Model-based effect estimation with compositional variables has a long history and was perhaps first thoroughly investigated under the name ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002). While parts of the literature focused on the problem of selecting design points $z_{1},\dots,z_{n}\in\Delta^{d-1}$ in optimal ways, a key aspect was also the problem of constructing reasonable regression models in this setting. Scheffé (1958) pointed out that polynomial models are overidentified with compositional data due to the sum-to-one constraint and provided a canonical choice of polynomials. Becker (1968) critiqued the use of polynomials for compositional features and instead provided parametric models with homogenous functions of degree one as building blocks. Cox (1971) argued that the polynomials of Scheffé would often not result in interpretable coefficients but that one should instead define the coefficients differently, e.g. so that the $j$th coefficient in a linear model corresponds to increasing the value of $Z^{j}$ while scaling the remaining components towards a fixed and problem-dependent ‘standard mixture’ $z_{*}\in\Delta^{d-1}$. It is possible to define a model-agnostic version of Cox’s linear effects using our perturbation-based framework (see Proposition S4 in Section S6 of the supplementary material). Draper and John (1977) developed predictive polynomial models based on the reciprocal of the components of $Z$ to cope with cases where the expected response diverges at the boundary. Aitchison and Bacon-Shone (1984) introduced the log-contrast regression model where the linear version corresponds to a linear regression on the log-transformed $Z$ with a sum-to-zero constraint on the coefficients. Log-contrast models have seen extensive use in practice and have since been generalized to high-dimensional $Z$ (Lin et al., 2014; Simpson et al., 2021). One of our proposed feature influence measures, the compositional feature influence with multiplicative speed (see Example 3), corresponds exactly to the log-contrast coefficients when the model is correctly specified (see Proposition S3 in Section S6 of the supplementary material or Huang et al. (2023) where this was first shown). Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022).
@@ 1.1 Related literature
Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis. It is well-known that many conventional statistical procedures become invalid when applied to compositional data. For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint. The most prevalent class of general methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach. In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982, 1986; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space. Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3). The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however this in principle requires $Z$ to take values in the open simplex. This is known as the zero problem and has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach. Greenacre et al. (2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches. While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022), or by appealing to more general notions of object-oriented data analysis (Marron and Dryden, 2021). There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022).
@@ 2 Target parameters via perturbations
    $v_{\psi}(z):=\mathds{1}_{\{\omega_{\psi}(z)\neq 0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
@@
Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex (see Section S1 of the supplementary material for a precise definition) at $z$, we define the average (directional) perturbation effect by
@@
We showcase different potential applications for both types of perturbation effect in the numerical experiments in Section 4. For binary perturbations, the denominator is included to scale the effect by the probability that the perturbation does not change $Z$. This is not required but in many cases provides a more meaningful parameter if one is interested in how the perturbation affects points on the simplex that are not yet at their endpoint. For instance, the denominator makes $\lambda_{\psi}$ invariant to modifications of the distribution of $Z$ that only changes $\mathbb{P}(Z\neq\psi(Z,1))$, i.e. the probability that $Z$ is not already perturbed. While both target parameters have obvious similarities to causal effects, they are a priori only (observational) statistical quantities. In particular, they are identifiable from the observational distribution as long as the conditions in Definition 2 are satisfied. We discuss the assumptions required for causal conclusions in Section 2.3.
Direct estimation of both target parameters relies on estimating the regression function $f:z\mapsto\mathbb{E}[Y\mid Z=z]$ and taking averages (see Section 3). An alternative estimation procedure for the effect of a binary perturbation $\lambda_{\psi}$ can be derived from noting that, if $L=\mathds{1}_{\{Z=\psi(Z,1)\}}$ and $W=\psi(Z,1)$, then
$\mathbb{E}[Y\,|\,L=1,W]=\frac{\mathbb{E}[YL\,|\,W]}{\mathbb{E}[L\,|\,W]}=\frac{\mathbb{E}[f(Z)L\,|\,W]}{\mathbb{E}[L\,|\,W]}=f(W).$
@@ 2.1 Perturbations on the simplex
We now turn to the construction of perturbations on the simplex. For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1). For a fixed endpoint there is a canonical (potentially geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint. The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation. We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from an interpretable summary statistic, which decreases when moving away from the endpoint, and interpreting the resulting effect as the effect on the response of slightly increasing the value of the statistic. For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
@@ 2.1.1 Endpoints in the simplex
As formally defined in Definition 1, an endpoint function $E:\Delta^{d-1}\rightarrow\Delta^{d-1}$, if it exists, determines where the perturbation ends when $\gamma$ is increased as much as possible. The simplest type of endpoint function on the simplex is a constant function, that is, there exists a fixed point $z_{*}\in\Delta^{d-1}$ such that for all $z\in\Delta^{d-1}$ it holds that $E(z)=z_{*}$. For example, we can choose the fixed endpoint $z_{*}=e_{j}$, which corresponds to a vertex on the simplex and hence a perturbation with this endpoint moves proportions from the other coordinates into the $j$th one. A further example, is to use the center of the simplex, i.e. $z_{*}=z_{\mathrm{cen}}$. As alluded to in the introduction, perturbations with $z_{\mathrm{cen}}$ as the endpoint can be seen as perturbing observations in the direction of equal distribution and hence capture a notion of diversity. For later reference, we call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto z_{\mathrm{cen}}$ centering perturbations.
A second type of endpoint function, that can be non-constant, is given by mapping each $z\in\Delta^{d-1}$ to amalgamations on the simplex. An amalgamation is a well-studied operation (e.g., Aitchison, 1986) which consist of merging parts of the compositional components into others. Formally, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint. Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed. In Section 2.1.3 we describe how working in other geometries can lead to different notions of direction. Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect. We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, where the effect of interest corresponds to increasing the value of the statistic, and then constructing a corresponding speed. In both cases, the construction of a derivative-isolating reparametrization is similar.
@@
Intuitively, a derivative-isolating reparametrization for $\psi$ should isolate changes along the direction $v_{E}$ from changes due to varying directions. We therefore consider the reparametrization $\phi_{\mathrm{unit}}:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(-\|E(z)-z\|_{1},(E(z),v_{E}(z)))$, where $\mathcal{W}$ is formally defined in Proposition 2. The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=\ell w_{v}+w_{E}$, so
@@
and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$. Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction to start from either a given interpretable speed or a summary statistic. The proof of the following result is given in Section S2.2 of the supplementary material.
@@
For $\mathcal{Z}\subseteq\Delta^{d-1}$, let $E:\mathcal{Z}\to\Delta^{d-1}$ be an endpoint function satisfying for all $z\in\mathcal{Z}$ that $E(z)\neq z$. Denote by $\mathcal{E}:=E(\mathcal{Z})$ the set of endpoints and define for all $z^{*}\in\mathcal{E}$ and all $u\in\mathbb{R}^{d}$ the sets $\mathcal{I}_{(z^{*},u)}:=\{\delta\in(0,\infty)\mid z^{*}+\delta u\in\Delta^{d-1}\}$ and $\mathcal{W}:=\{(w_{E},w_{v})\in\mathcal{E}\times\mathbb{R}^{d}\mid\|w_{v}\|_{1}=1,\,\mathcal{I}_{w}\neq\emptyset\}$. Finally, define the unit speed reparametrization $\phi_{\mathrm{unit}}=(\phi_{\mathrm{unit}}^{L},\phi_{\mathrm{unit}}^{W}):\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}^{L}(z):=-\|E(z)-z\|_{1}$ and $\phi_{\mathrm{unit}}^{W}(z):=(E(z),v_{E}(z))$ with $v_{E}$ defined in (6). Then the following approaches lead to perturbations with corresponding derivative-isolating reparametrizations.
@@
    (Speed) Let $s:\mathcal{Z}\rightarrow(0,\infty)$ be a given speed function. For all $w\in\mathcal{W}$ define $s_{w}:\mathcal{I}_{w}\rightarrow(0,\infty)$ for all $\delta\in\mathcal{I}_{w}$ by $s_{w}(\delta):=s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and let $t_{w}:\mathcal{I}_{w}\to\mathbb{R}$ denote a solution to
@@
    Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t_{\phi_{\mathrm{unit}}^{W}(z)}(\|E(z)-z\|_{1}),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$ and any other perturbation with the same speed and direction.
@@
    (Summary statistic) Let $t:\mathcal{Z}\rightarrow\mathbb{R}$ be a given summary statistic. For all $w\in\mathcal{W}$ define $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ for all $\delta\in\mathcal{I}_{w}$ by $t_{w}(\delta):=t(\phi_{\mathrm{unit}}^{-1}(-\delta,w))$ and assume that $t_{w}$ is differentiable and strictly decreasing. Then, the reparametrization $\phi:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi(z):=(t(z),\phi_{\mathrm{unit}}^{W}(z))$ is derivative-isolating for the perturbation $\psi:z\mapsto z+\gamma s(z)v_{E}(z)$, where the speed $s:\mathcal{Z}\rightarrow\mathbb{R}$ is defined for all $z\in\mathcal{Z}$ as
@@
Suppose we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in\Delta^{d-1}$. We can investigate the effect of such a change by means of a $([d]\setminus\{j\}\to\{j\})$-amalgamating perturbation, that is, a perturbation which satisfies for all $z\in\Delta^{d-1}\setminus\{e_{j}\}$ that $v_{\psi}(z)=\frac{e_{j}-z}{\|e_{j}-z\|_{1}}$. While for some purposes it may be perfectly reasonable to use a constant speed $s_{\psi}\equiv 1$, in other settings this speed can be unintuitive. For example, if we imagine the observed $Z$ as being generated from an unobserved $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale by $Z:=\mathsf{C}(X)=\frac{1}{\|X\|_{1}}X$, it might be more natural to consider which speeds are induced on the simplex from hypothetical changes to $X^{j}$. A first attempt might be to consider the speed resulting from adding a constant $c$ to $X^{j}$. However, since
@@
this speed is not scale-invariant for $X$, that is, the speed when considering the counts $X$ and $cX$ for $c>0$ may differ. This means that the speed on the simplex would depend on the unobserved quantity $\|X\|_{1}$, so the perturbation is ill-defined. Instead, we can consider the speed resulting from multiplying $X^{j}$ by $1+c$ (so that $c=0$ corresponds to no change) which yields a well-defined perturbation on the simplex since this operation is scale-invariant. More formally, letting $\odot$ denote the Hadamard (point-wise) product, the resulting speed is
@@
which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition 2 (a). Since for all $w\in\mathcal{W}$ it holds that $w_{E}^{j}=1$ (component $w_{E}$ corresponds to the endpoint $E(z)$) and $w_{v}^{j}=\frac{1}{2}$ (component $w_{v}$ corresponds to the direction $v_{E}(z)$ and we used that $v_{E}(z)^{j}=\frac{1-z^{j}}{\|e_{j}-z^{j}\|_{1}}=\frac{1-z^{j}}{2(1-z^{j})}=\frac{1}{2}$), we have that $\phi_{\mathrm{unit}}^{-1}(\ell,w)^{j}=\frac{\ell}{2}+1$. Therefore,
$s_{w}:\delta\mapsto s(\phi_{\mathrm{unit}}^{-1}(-\delta,w))=\frac{(2-\delta)\delta}{2}.$
@@
$\phi^{W}(z)=(e_{1},v_{E}(z))\quad\text{and}\quad\phi^{L}(z)=t_{\phi^{W}(z)}(\|e_{j}-z\|_{1})=\log\left(\frac{z^{j}}{1-z^{j}}\right)$